# Hey Dashboard!: Supporting Voice, Text, and Pointing Modalities in Dashboard Onboarding

### v1

@@ 2.4 Automatic Generation of Dashboards
The complexity of creating effective dashboards has motivated significant research into automated generation approaches. Early systems such as VizDeck [21] introduced the concept of self-organizing dashboards that automatically recommend visualizations based on statistical properties of data, using a card game metaphor to help users organize visualizations into interactive dashboards with minimal programming effort.
@@
Deep learning approaches have shown promise for dashboard recommendation and generation. Wu et al. [46] developed MultiVision, which uses deep learning models to assist in analytical dashboard design by learning from provenance data and authoring logs. Their mixed-initiative system allows users to provide optional input while the model recommends data column selections and multiple chart combinations. Similarly, Deng et al. [11] proposed DashBot, which uses deep reinforcement learning to generate insight-driven dashboards by constructing training environments based on visualization knowledge.
@@
Intent-based recommendation systems have emerged to support dashboard composition workflows. Elshehaly et al. [17] developed QualDash for healthcare quality improvement, introducing a metric card metaphor that serves as a building block for generating adaptable dashboards across different hospital units. Pandey et al.  [31] introduced MEDLEY, which recommends dashboard collections based on analytical intents such as measure analysis, change analysis, and distribution analysis. Users can specify intents explicitly or implicitly through data attribute selection, with the system providing collections of logically grouped views and filtering widgets.
@@
The proliferation of large language models has enabled new approaches to dashboard creation through natural language interaction. Dibia [14] developed LIDA, which uses LLMs for automatic generation of grammar-agnostic visualizations and infographics, demonstrating how conversational interfaces can simplify visualization creation. Building on this foundation, Lisnic et al. [24] developed a LLM-powered system specifically for creating text content within dashboards, showing how artificial intelligence (AI) can assist with narrative elements that provide context and explanation.
@@
Recent work has integrated LLMs into broader visual analytics workflows. Chen et al. [8] developed InterChat, which combines direct manipulation with natural language interaction for visual analytics, including LLM-based prompting capabilities. Zhao et al. [48] proposed LightVA, a framework that uses LLM agents for task planning and execution in visual analytics, employing a recursive process with planner, executor, and controller components. Wen et al. [45] explored multimodal prompting with LLMs for visualization authoring, investigating how different input modalities can enhance the creation process beyond traditional text-based interaction.
@@
While these automated generation approaches have made dashboard creation more accessible, most focus on the authoring process rather than helping users understand and navigate existing dashboards. The challenge of onboarding users to complex, pre-existing dashboard systems—particularly through multimodal interaction—remains largely unaddressed by current automatic generation methods.
3 Diana: Dashboard Interactive Assistant for Navigation and Analysis
@@
Figure 2: Diana User Interface. The left side shows a visualization dashboard with Diana activated: (a) represents the lasso selection button, (b) push-to-talk for speech-based interaction, and (c) chat-based interaction. Currently, the lasso selection is activated and the y-axis of the visualization is highlighted, along with the radial menu which is opened for the y-axis information. The right side shows multimodal interaction with mouse-based lasso selection and chat-based interaction that infers the context of the conversation based on the selected visual.
@@
Existing onboarding methods for dashboards, such as documentation, training, videos, and meetings, often fail to support users in situ. They are static, linear, and cannot adapt to individual preferences or questions that may arise during the onboarding process. Additionally, the user needs to constantly switch programs, manually test interactions while cross-referencing instructions, which can lead to frustration with onboarding help. Self-designed tutorials, such as step-by-step interfaces, also face the challenge of being ignored by many users. To address the balance between autonomy and agency, we designed Diana (Dashboard Interactive Assistant for Navigation and Analysis), a multimodal onboarding assistant powered by Large Language Models.
@@
Diana is designed not to replace domain expertise, but to bridge the gap between static onboarding and in-situ analysis. Its role is to help users navigate dashboards, understand visualizations, and learn how to perform analysis through guided, multimodal interaction. Users can interact with Diana via three interaction modalities: keyboard (chat), mouse (lasso selection), and voice (speech).
Based on the chosen modality, Diana provides feedback through text, audio, a contextual menu, and visual highlights to orient users within the dashboard.
@@
We designed Diana according to these goals that support the user’s orientation and onboarding within the dashboard. It is implemented using a Microsoft Power BI dashboard embedded in a custom web application, thereby satisfying DG1. We distinguish between modalities, which define how users interact with Diana (keyboard, mouse, voice), and features, which define how the system provides feedback and orientation (contextual help menu, visual highlighting). We explain them below.
3.1 Mouse-based Interaction
@@
The mouse is the primary input device in most dashboard environments, and users are generally familiar with interacting with visualizations through clicking. To extend this familiar mode of interaction for onboarding (DG2), we included two complementary features: lasso selection (or free-form selection) and hover functionality. Lasso selection enables users to mark a region of interest within the dashboard, mimicking the pointing gestures a human expert might use when directing attention to a particular element. Once a region is selected, the system captures both the highlighted area and the corresponding visualization. At the same time, the dashboard remains fully interactive for further exploration.
@@
Hover functionality provides additional detail in two ways: (i) through contextual help options in the radial menu (see Section 3.4), and (ii) through embedded hyperlinks returned in responses to text-based queries (see Section 3.3). This way lasso serves as a way of anchoring questions to a region, while hover enriches that region with further information. In a typical onboarding scenario, users also point and ask questions, therefore, through mouse interaction, we provide human-like onboarding support (DG3) and also facilitate orientation by anchoring onboarding interactions to specific areas of the visualization (DG4).
3.2 Voice-based Interaction
@@
With the growing support for multimodal interaction in LLM-powered assistants (e.g., Microsoft Co-Pilot), we included a voice modality to mimic the experience of conversing with a human onboarder (DG2 and DG3). Speech input allows users to issue queries in natural language and receive real-time guidance without relying on typing.
@@
Users can enable this modality via a push-to-talk model, where the users can press and hold the mouse button while speaking, and upon release, Diana processes the input using the Realtime API [2] and immediately returns a spoken response (DG5). This design ensures that users maintain full control over when the system listens, avoiding interruptions during regular conversation. We initially experimented with a wake-word approach (“Hey dashboard!”), but found it prone to false activations, as the system occasionally responded to background discussions that were not intended for the assistant. We therefore selected the push-to-talk model as a more reliable and less intrusive interaction style. In addition to the audio output, Diana also highlights specific regions of the dashboard to direct user’s attention (DG4).
3.3 Keyboard-based Interaction
@@
The keyboard-based interaction complements the voice-based interaction by offering a written chat interface that supports deliberate, text-based interaction. Unlike spoken dialogue, written exchanges create a persistent trace of the conversation, allowing users to revisit instructions and follow them step by step (DG2 and DG3).
@@
The chat component uses the OpenAI APIs to process user queries and provides responses in textual form. In addition to plain text, these responses may also include embedded hyperlinks and hovering over them highlights the corresponding regions on the dashboard (e.g., the visualization, its legends, or its axes) (DG4 and DG5).
3.4 Contextual Help Menu
@@
To support onboarding in a deterministic and structured way, we designed a contextual radial menu that complements the natural language modalities (chat and speech), which rely heavily on LLMs for generating responses. While conversational interaction adapts to the user’s phrasing and provides guidance step by step, the menu makes the full range of onboarding information visible, even if the user does not know what to ask.
@@
We extended the component graph for dashboards described by Dhanoa et al. [12] for the design of our contextual help menu. As a visualization can contain many elements, such as data, legends, axes, and interaction options, that may require explanation, we adopted a bottom-up approach to assemble this information hierarchically. To present this information visually, we used a circular radial layout that progressively reveals detail across levels. It captures the complexity of the component graph while presenting it in a structured and navigable form.
@@
The menu appears automatically when a region of a chart is selected with the lasso (DG4). It opens at the relevant category and displays a pop-up information text box with a description of the selected region. Users can also hover over menu items, prompting the text box to display the corresponding explanation.
Although the schema is deterministic, we used an LLM to polish the narrative text associated with each menu entry. For instance, instead of a bare label such as Currency Scale, the menu might display The X-axis represents values in USD on a continuous scale. In this way, the LLM enriches the text with more natural explanations, while the structure and content remain grounded in the underlying visualization metadata. In addition to plain text, icons are used to enhance readability.
@@
3.5 Visual Highlighting
We designed the visual highlighting feature to complement all three modalities by directing user attention to specific dashboard elements, similar to how a human onboarder might point to a chart or annotate it (DG3). When triggered through a query, whether via lasso selection, voice command, or a written input, the system highlights the relevant visualization, its axes, or its legend. These highlights appear as rectangular pulsating bars that serve as visual cues, helping users orient themselves and understand where to focus (DG4).
@@ 3.5 Visual Highlighting
3.6 Combining Modalities and Features
@@
In real onboarding scenarios, users rarely rely on a single modality when seeking onboarding support. Therefore, we designed Diana to allow users to combine modalities and features, in parallel or sequentially, to reproduce the onboarding experience closer to that provided by a human onboarder (DG3). Our system enables users to lasso a region (mouse-based interaction) and simply use the voice-based interaction to ask a question for the selected area. The user can also view more detailed information in the contextual menu, while receiving audio onboarding. The use of visual highlights can prompt further exploration, prompting the user to open the contextual menu on that specific visualization or simply selecting that visualization and asking a chat-based query (keyboard-interaction).
@@
These combinations allow Diana to have different kinds of support: the mouse-based anchoring to a selected region, the contextual menu for structured information, natural language support in the form of voice and keyboard-based interactions, and visual highlighting for orientation and directed attention. Altogether, they can create an adaptive onboarding experience in which onboarding can be tailored to both the task at hand and the user’s preferred way of interacting.
@@
Unlike typical LLM-based assistants that aim to deliver direct answers, Diana focuses on orientation and learning. The system is prompted with information about the structure of the dashboard (such as available visuals, axes, maps, and overview descriptions) but not with the underlying data values. As a result, Diana cannot return analytical results itself, except for basic descriptive queries (e.g., what is the y-axis?). This design choice ensures that the assistant focuses on orientation and guidance rather than replacing the user’s own analytical work. We explain the working of the system in the next section.
4 Implementation
We implemented Diana using the Microsoft Power BI Embedded report [27], to integrate our solution within existing commercial dashboard tool, such as Microsoft Power BI. We obtain information about all the visualizations in the dashboard using Power BI’s REST API [35]. We also apply simple heuristics to deduce the information not present within these APIs. Additionally, we use OpenAI chat completion models [1] and Realtime API [2] for chat and voice-based interactions, respectively.
The system itself is built using the Vue framework [3] for the front-end and Flask API for the Python backend. All calls to the Open AI APIs are routed through the backend. We also created prompts for chat and voice-based interactions; they are provided in the supplementary materials. These prompts are designed such that the system does not output answers, but rather define steps for getting to an answer. The reason for this is that the purpose of Diana is not to create yet another assistant that solves user queries, but that provides onboarding and navigation support so that users can remain responsible for the analytical tasks.
We implemented Diana using Microsoft Power BI Embedded [27] to integrate within a widely used commercial dashboard platform without building a custom dashboard stack. We obtain visualization information, such as titles, axes, data, and layout bounds via the Power BI Embed API for Typescript [25], and use the Power BI REST API [35] for embedding logistics (e.g., report loading and issuing logging in token). When certain attributes are unavailable, we apply lightweight heuristics to infer them from the rendered view.

### v2

@@ 2.4. Automatic Generation of Dashboards
The complexity of creating effective dashboards has motivated significant research into automated generation approaches. Early systems such as VizDeck (Key et al., 2012) introduced the concept of self-organizing dashboards that automatically recommend visualizations based on statistical properties of data, using a card game metaphor to help users organize visualizations into interactive dashboards with minimal programming effort.
@@
Deep learning approaches have shown promise for dashboard recommendation and generation. Wu et al. (Wu et al., 2022) developed MultiVision, which uses deep learning models to assist in analytical dashboard design by learning from provenance data and authoring logs. Their mixed-initiative system allows users to provide optional input while the model recommends data column selections and multiple chart combinations. Similarly, Deng et al. (Deng et al., 2023) proposed DashBot, which uses deep reinforcement learning to generate insight-driven dashboards by constructing training environments based on visualization knowledge.
@@
Intent-based recommendation systems have emerged to support dashboard composition workflows. Elshehaly et al. (Elshehaly et al., 2021) developed QualDash for healthcare quality improvement, introducing a metric card metaphor that serves as a building block for generating adaptable dashboards across different hospital units. Pandey et al.  (Pandey et al., 2022) introduced MEDLEY, which recommends dashboard collections based on analytical intents such as measure analysis, change analysis, and distribution analysis. Users can specify intents explicitly or implicitly through data attribute selection, with the system providing collections of logically grouped views and filtering widgets.
@@
The proliferation of large language models has enabled new approaches to dashboard creation through natural language interaction. Dibia (Dibia, 2023) developed LIDA, which uses LLMs for automatic generation of grammar-agnostic visualizations and infographics, demonstrating how conversational interfaces can simplify visualization creation. Building on this foundation, Lisnic et al. (Lisnic et al., 2025) developed a LLM-powered system specifically for creating text content within dashboards, showing how artificial intelligence (AI) can assist with narrative elements that provide context and explanation.
@@
Recent work has integrated LLMs into broader visual analytics workflows. Chen et al. (Chen et al., 2025) developed InterChat, which combines direct manipulation with natural language interaction for visual analytics, including LLM-based prompting capabilities. Zhao et al. (Zhao et al., 2025) proposed LightVA, a framework that uses LLM agents for task planning and execution in visual analytics, employing a recursive process with planner, executor, and controller components. Wen et al. (Wen et al., 2025) explored multimodal prompting with LLMs for visualization authoring, investigating how different input modalities can enhance the creation process beyond text-based interaction.
@@
While these automated generation approaches have made dashboard creation more accessible, most focus on the authoring process rather than helping users understand and navigate existing dashboards. The challenge of onboarding users to complex, pre-existing dashboard systems—particularly through multimodal interaction—remains largely unaddressed by current automatic methods.
3. Diana: Dashboard Interactive Assistant for Navigation and Analysis
@@
Figure 2. Diana user interface. The first image shows a visualization dashboard with Diana activated: (a) represents the lasso selection button, (b) push-to-talk for speech-based interaction, and (c) chat-based interaction. Currently, the lasso selection is activated and the y-axis of the visualization is highlighted, along with the radial menu which is opened for the y-axis information. The second image shows multimodal interaction with mouse-based lasso selection and chat-based interaction that infers the context of the conversation based on the selected visual.
The interface displays key revenue metrics along the top, followed by a funnel chart, a monthly line chart, a horizontal bar chart of revenue by product category, and a world map showing geographic revenue distribution. A tooltip explains the sorting of the vertical axis in the bar chart. The pattern shown across the dashboard is that the revenue goal exceeds the revenue won by a significant margin, the funnel chart demonstrates a typical narrowing from Proposal to Closed stages, and the monthly line chart suggests seasonal variability with a noticeable rise late in the year. The bar chart indicates that Software and Services are the strongest revenue categories, while Accessories generate the least. The world map highlights North America and Europe as the most revenue-dense regions, reinforcing the concentration of performance in specific geographic areas. The second figure has the emphasis on guided interpretation: the help panel clarifies how to read the map and highlights the contrast between regions, demonstrating how system-generated explanations scaffold the user’s understanding of broader revenue patterns.
@@
Existing onboarding methods for dashboards, such as documentation, training, videos, and meetings, often fail to support users in situ. They are static, linear, and cannot adapt to individual preferences or questions that may arise during the onboarding process. Additionally, the user needs to constantly switch programs, and manually test interactions while cross-referencing instructions, which can lead to frustration with onboarding help. Self-designed tutorials, such as step-by-step interfaces, also face the challenge of being ignored by many users. To address the balance between autonomy and agency, we designed Diana (Dashboard Interactive Assistant for Navigation and Analysis), a multimodal onboarding assistant powered by Large Language Models.
@@
Diana is designed not to replace domain expertise, but to bridge the gap between static onboarding and in-situ analysis. Its role is to help users navigate dashboards, understand visualizations, and learn how to perform analysis through guided, multimodal interaction. Users can interact with Diana via three interaction modalities: keyboard (chat), mouse (lasso selection), and voice (speech). Based on the chosen modality, Diana provides feedback through text, audio, a contextual menu, and visual highlights to orient users.
@@
We designed Diana according to these goals that support the user’s orientation and onboarding within the dashboard. It is implemented using a Microsoft Power BI dashboard embedded in a custom web application, thereby satisfying DG1 (familiar). We distinguish between modalities, which define how users interact with Diana (keyboard, mouse, voice), and features, which define how the system provides feedback and orientation (contextual help menu, visual highlighting). We explain them below.
3.1. Mouse-based Interaction
@@
The mouse is the primary input device in most dashboard environments, and users are generally familiar with interacting with visualizations through clicking. To extend this familiar mode of interaction for onboarding (DG2 (multimodal)), we included two complementary features: lasso selection (or free-form selection) and hover functionality (see Figure 2 (a)). Lasso selection enables users to mark a region of interest within the dashboard, mimicking the pointing gestures a human expert might use when directing attention to a particular element. Once a region is selected, the system captures the rectangular coordinates of the highlighted area and maps them to the corresponding visualizations. At the same time, the dashboard remains fully interactive for further exploration.
@@
Hover functionality provides additional detail in two ways: (i) through contextual help options in the radial menu (see Section 3.4), and (ii) through embedded hyperlinks returned in responses to text-based queries (see Section 3.3). Therefore, lasso serves as a way of anchoring questions to a region, while hover enriches that region with further information. In a typical onboarding scenario, users also point and ask questions. Mouse-based interactions allow us to provide human-like onboarding support (DG3 (human-like)) and also facilitate orientation by anchoring onboarding interactions to specific areas of the visualization (DG4 (orientation)).
3.2. Voice-based Interaction
@@
With the growing support for multimodal interaction in LLM-powered assistants (e.g., Microsoft Copilot), we included a voice modality to mimic the experience of conversing with a human onboarder (DG2 (multimodal) and DG3 (human-like)). Speech input allows users to issue queries in natural language and receive real-time guidance without relying on typing.
@@
Users can enable this modality via a push-to-talk model(see Figure 2 (b)), where they press and hold the mouse button while speaking, and upon release, Diana processes the input using the Realtime API (OpenAI, 2023b) and immediately returns a spoken response (DG5 (guidance & autonomy)). This design ensures that users maintain full control over when the system listens, avoiding interruptions during regular conversation. We initially experimented with a wake-word approach (“Hey dashboard!”), but found it prone to false activations, as the system occasionally responded to background discussions that were not intended for the assistant. We therefore selected the push-to-talk model as a more reliable and less intrusive interaction style. In addition to the audio output, Diana also highlights specific regions of the dashboard to direct the user’s attention (DG4 (orientation)).
3.3. Keyboard-based Interaction
@@
The keyboard-based interaction(see Figure 2 (c)) complements the voice-based interaction by offering a written chat interface that supports deliberate, text-based interaction. Unlike spoken dialogue, written exchanges create a persistent trace of the conversation, allowing users to revisit instructions and follow them step by step (DG2 (multimodal) and DG3 (human-like)).
@@
The chat component uses the OpenAI APIs to process user queries and returns responses in text. In addition to plain text, these responses may also include embedded hyperlinks, and hovering over these responses highlights the corresponding regions on the dashboard (e.g., the visualization, its legends, or its axes) (DG4 (orientation) and DG5 (guidance & autonomy)).
3.4. Contextual Help Menu
@@
To support onboarding in a deterministic and structured way, we designed a contextual radial menu that complements the natural language modalities (chat and speech), which rely heavily on LLMs for generating responses. While conversational interaction adapts to the user’s phrasing and provides guidance step by step, the menu makes the full range of onboarding information visible, even if the user does not know what to ask (Ko and Myers, 2004).
@@
We extended the component graph for dashboards described by Dhanoa et al. (Dhanoa et al., 2022) for the design of our contextual help menu. As a visualization can contain many elements, such as data, legends, axes, and interaction options, that may require explanation, we adopted a bottom-up approach to assemble this information hierarchically. To present this information visually, we used a circular radial layout (CRL) that progressively reveals detail across levels. Pourmemar et al. (Pourmemar and Poullis, 2019) found that using the CRL was faster than the hierarchical drop-down menu in an immersive augmented reality environment. Samp and Decker also found that pointing is faster with CRL and it is beneficial for the design of displayed hierarchical menus (Samp and Decker, 2010).
Therefore, we also used CRL to capture the complexity of the component graph in a structured and navigable form.
@@
The menu appears automatically when a region of a chart is selected with the lasso (DG4 (orientation)). It opens at the relevant category and displays a pop-up information box with a description of the selected region.
When users hover over menu items, the information box displays the corresponding explanation.
Although the schema is deterministic, we used an LLM to polish the narrative text associated with each menu entry. For instance, instead of a bare label such as Currency Scale, the menu might display The X-axis represents values in USD on a continuous scale. In this way, the LLM enriches the text with more natural explanations, while the structure and content remain grounded in the underlying visualization metadata. In addition to plain text, we use icons to enhance readability.
@@
3.5. Visual Highlighting
We designed the visual highlighting feature to complement all three modalities by directing user attention to specific dashboard elements, similar to how a human onboarder might point to a chart or annotate it (DG3 (human-like)).
When a user lasso-selects a region, the system maps its rectangular coordinates to identify which visualization components they intersect. For keyboard or voice-based interactions, it extracts keywords (e.g., title, axis, legend) and maps them to the corresponding visualization components based on words in the user’s input (written or spoken).
These highlights appear as pulsating rectangular bars that serve as visual cues, helping users orient themselves and understand where to focus (DG4 (orientation)).
@@ 3.5. Visual Highlighting
3.6. Combining Modalities and Features
@@
In real onboarding scenarios, users rarely rely on a single modality when seeking onboarding support. Therefore, we designed Diana to allow users to combine modalities and features, in parallel or sequentially, to reproduce an onboarding experience similar to that provided by a human onboarder (DG3 (human-like)). Our system enables users to lasso a region (mouse-based interaction) and simply use the voice-based interaction to ask a question for the selected area. The user can also view more detailed information in the contextual menu, while receiving audio onboarding. The use of visual highlights can prompt further exploration, leading the user to open the contextual menu for that visualization or simply select it and ask a chat-based query (keyboard interaction).
@@
These combinations allow Diana to provide different kinds of support: mouse-based anchoring to a selected region, a contextual menu for structured information, natural language support via voice and keyboard-based interactions, and visual highlighting for orientation and directed attention. Altogether, they can create an adaptive experience in which onboarding can be tailored to both the task at hand and the user’s preferred way of interacting.
@@
Unlike typical LLM-based assistants that aim to deliver direct answers, Diana focuses on orientation and learning. The system is prompted with information about the dashboard’s structure (such as available visuals, axes, maps, and overview descriptions) but not with the underlying data values. As a result, Diana cannot return analytical results itself, except for basic descriptive queries (e.g., what is the y-axis?). This design choice ensures that the assistant focuses on orientation and guidance rather than replacing the user’s own analytical work. We next explain how Diana works.
4. Implementation
We implemented Diana using the Microsoft Power BI Embedded report (Microsoft Corporation, 2023a) to integrate our solution within an existing commercial dashboard tool such as Microsoft Power BI. We obtain information about all visualizations in the dashboard using Power BI’s REST API (Sharabi et al., 2023). We also apply simple heuristics to deduce information not exposed through these APIs, such as determining the positions of legends and axes, which are not provided by default. We infer these by adding padding offsets to the top-left rectangular coordinates of each visualization based on its type. This allows us to construct an internal map of the dashboard layout and the relative placement of its components. As different visualization types place their axes in different locations, we extract those positions as well by checking them against specific Power BI keywords. We also create a transparent overlay on the dashboard for lasso selections. When a user draws a region, we capture its bounding box and map it onto the dashboard layout. The system itself is built using the Vue framework (You, 2023) for the frontend and Flask for the Python backend.
Using Vue, we implement dedicated components for all modalities and features, including visual highlights and the radial menu. When a user selects a region, the system captures its coordinates and maps them onto the dashboard layout to answer pointed questions. It then activates the appropriate Vue components, such as those for visual highlighting and the radial menu, which are rendered as overlay layers on top of the embedded Power BI iFrame.
When the user interacts through another modality, such as the keyboard or speech, the system forwards the relevant visualization metadata to the backend. The backend then integrates this contextual information with our prompts to generate appropriate onboarding instructions. The system maintains this context until the user selects a different region or asks a new question.